OpenStudy (anonymous):

At the Wild Thing Zoo, you can rent a motorized cart to tour the grounds for a $4 initial charge and $4 per hour. An equation that represents the total cost y for x hours is y = 4 + 4x. At Safari Zoo, you can rent the same cart for a $3 inital charge and $5 for hour. An equation that represents the total cost y for x hours is y = 3 + 5x. For how many hours is the total charge the same?

OpenStudy (anonymous):

Some 1 Help Please!!!!!

OpenStudy (radar):

y=y 4 + 4x = 3 + 5x 1=x x=1

OpenStudy (radar):

Did you follow with understanding?